OpenStudy (anonymous):

Cubed root of -25 * cubed root of 5

Nnesha (nnesha):

you can convert radical to an exponent form \[\huge\rm \sqrt[\color{Red}{n}]{x^\color{blue}{m}}=x^\frac{\color{blue}{ m }}{ \color{ReD}{n} }\]index becomes denominator of the fraction

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \sqrt[3]{-25}\times \sqrt[3]{5}=\sqrt[3]{-25\cdot 5}=\sqrt[3]{-125} }\) And then: \(\large\color{black}{ \displaystyle (-5)^3=-125}\)

Nnesha (nnesha):

\[\huge\rm \sqrt[3]{-25} *\sqrt[3]{5}\] or since the both are cube root you can multiply the numbers

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha you can convert radical to an exponent form \[\huge\rm \sqrt[\color{Red}{n}]{x^\color{blue}{m}}=x^\frac{\color{blue}{ m }}{ \color{ReD}{n} }\]index becomes denominator of the fraction \(\color{blue}{\text{End of Quote}}\) for this one you have to rewrite 25 in terms of base 5